# Q-BAR: Blogger Anomaly Recognition via Quantum-enhanced Manifold Learning

### v1

@@ 4.2 Test Set Generation (Simulation of Malice)
Since ground-truth “malicious edits” are difficult to harvest at scale with reliable labels, we simulated the behaviour of marketing accounts to create a robust test set. We selected 20 bloggers from the pool and generated test samples using two distinct adversarial strategies. The first strategy, Semantic Splicing (50% of test data), involved segmenting original videos and randomly reordering sentences or removing conditional clauses (e.g., transforming “If X, then Y” to “Y”) to simulate out-of-context editing. The second strategy, Fake Expansion (50% of test data), employed Large Language Models (LLMs) to generate plausible but factually fabricated statements in the blogger’s specific style. We then utilized Text-to-Speech (TTS) and Lip-Sync technologies to insert these segments, simulating high-level fabrication that retains visual fidelity.
@@ 4.4 Baselines
We evaluated our proposed method against standard baselines to benchmark performance. For unimodal comparisons, we employed a Text-only baseline using BERT embeddings with a One-Class SVM (OC-SVM), and an Audio-only baseline analyzing spectral features. We also included a Visual-only (VAD) baseline based on traditional motion anomaly detection. For multimodal deep learning comparisons, we utilized Deep Support Vector Data Description (Deep SVDD) (ruff2018deep), which is the closest classical conceptual counterpart to our method, training a neural network to map data into a hypersphere. Additionally, we compared against a standard multimodal Autoencoder (AE) that utilizes reconstruction error for anomaly scoring.
@@ 4.7 Parameter Efficiency of Q-BAR
A primary contribution of this work lies in the resource efficiency of the quantum-enhanced model. While Q-BAR matches and slightly exceeds the performance of heavy classical deep learning models like Deep SVDD, it does so with a fundamental difference in resource utilization. Classical Deep SVDD requires thousands of parameters to stabilize the high-dimensional multimodal features to avoid underfitting. In contrast, our quantum circuit, leveraging the PEQAD framework (wang2025parameter), relies on as few as 240 trainable parameters (or 16 in our most compressed ablation) to achieve similar convergence.
@@ 6 Conclusion
In this work, we presented Q-BAR, a pioneering framework for Blogger Anomaly Recognition designed to detect semantic mutations in short video content. Recognizing the data scarcity inherent in modeling individual creators, we integrated the PEQAD strategy to map multimodal features into a creator-specific semantic manifold. Our experiments demonstrate that while the detection of high-level logic violations remains a challenging task, our quantum-enhanced model achieves detection performance comparable to deep classical baselines (F1 $\approx$ 0.71) while utilizing orders of magnitude fewer parameters. This efficiency enables rapid, low-resource deployment, offering a practical solution for platform-scale regulation. Future work will focus on bridging the gap between simulation and real quantum hardware, improving robustness against adversarial audio attacks, and refining the mutation axis to better distinguish malicious intent from benign satire.

### v2

@@ 4.2 Test Set Generation (Simulation of Malice)
Since ground-truth “malicious edits” are difficult to harvest at scale with reliable labels, we simulated the behaviour of marketing accounts to create a robust test set. We selected 20 bloggers from the pool and generated test samples using two distinct adversarial strategies. The first strategy, Semantic Splicing (50% of test data), involved segmenting original videos and randomly reordering sentences or removing conditional clauses (e.g., transforming “If X, then Y” to “Y”) to simulate out-of-context editing. The second strategy, Fake Expansion (50% of test data), employed Large Language Models (LLMs) to generate plausible but factually fabricated statements in the blogger’s specific style. We then utilized TTS and Lip-Sync technologies to insert these segments, simulating high-level fabrication that retains visual fidelity.
@@ 4.4 Baselines
We evaluated our proposed method against standard baselines to benchmark performance. For unimodal comparisons, we employed a Text-only baseline using BERT embeddings with a One-Class SVM (OC-SVM), and an Audio-only baseline analyzing spectral features. We also included a Visual-only (VAD) baseline based on traditional motion anomaly detection. For multimodal deep learning comparisons, we utilized Deep SVDD (Ruff et al., 2018), which is the closest classical conceptual counterpart to our method, training a neural network to map data into a hypersphere. Additionally, we compared against a standard multimodal Autoencoder (AE) that utilizes reconstruction error for anomaly scoring.
@@ 4.7 Parameter Efficiency of Q-BAR
A primary contribution of this work lies in the resource efficiency of the quantum-enhanced model. While Q-BAR matches and slightly exceeds the performance of heavy classical deep learning models like Deep SVDD, it does so with a fundamental difference in resource utilization. Classical Deep SVDD requires thousands of parameters to stabilize the high-dimensional multimodal features to avoid underfitting. In contrast, our quantum circuit, leveraging the PEQAD framework (Wang et al., 2025a), relies on as few as 240 trainable parameters (or 16 in our most compressed ablation) to achieve similar convergence.
@@ 6 Conclusion
In this work, we presented Q-BAR, a framework for Blogger Anomaly Recognition designed to detect semantic mutations in short video content. Recognizing the data scarcity inherent in modeling individual creators, we integrated the PEQAD strategy to map multimodal features into a creator-specific semantic manifold. Our experiments demonstrate that while the detection of high-level logic violations remains a challenging task, our quantum-enhanced model achieves detection performance comparable to deep classical baselines (F1 $\approx$ 0.71) while utilizing orders of magnitude fewer parameters. This efficiency enables rapid, low-resource deployment, offering a practical solution for platform-scale regulation. Future work will focus on bridging the gap between simulation and real quantum hardware, improving robustness against adversarial audio attacks, and refining the mutation axis to better distinguish malicious intent from benign satire.
Acknowledgements
We thank the content creators whose publicly available videos were used in constructing the dataset for this study. Their contributions to the online media ecosystem made it possible to investigate creator-specific semantic consistency and anomaly detection in real-world multimedia environments. M.W. acknowledges financial support from the UCL Dean’s Prize for doctoral research. The authors also acknowledge the Institute for Mathematical and Intelligent Systems at Wuhan University for providing computational resources, including access to NVIDIA GPU infrastructure used in this work. This work benefited from open-source quantum software frameworks, including PennyLane and NVIDIA CUDA-Q. This research represents the independent work of the authors, and the conclusions expressed in this paper do not necessarily reflect the views of the supporting institutions or the developers of the software frameworks used in this study.